OpenStudy (anonymous):

the function f(x)=x^4-4x^2 has how many mins and maxs?

OpenStudy (tkhunny):

Possible or for sure?

OpenStudy (anonymous):

for sure

hero (hero):

Hint: Graph the function

hero (hero):

Alternatively, set f(x) = 0 0 = x^4 - 4x^2 0 = x^2(x^2 - 4)

hero (hero):

Solve for x from there